# Exploring Lattice Thermal Conductivity Models via Interpretable Deep Learning to Accelerate Materials Discovery

### v7

@@ I Introduction
Lattice Thermal Conductivity (LTC) is a critical physical parameter that quantifies a material’s ability to transfer heat through lattice vibrations. It has diverse applications in thermal management [1], energy conversion [2], and thermoelectric materials [3]. High-LTC materials are used in electronic devices to efficiently dissipate heat and prevent overheating [4]. In contrast, low-LTC materials exhibit excellent performance in thermoelectric conversion, making them ideal for developing efficient thermoelectric generators (TEGs) [5] and coolers [6]. LTC is crucial for material design and optimization, yet acquiring it for specific materials is challenging.
@@
Traditional experimental methods for measuring LTC, such as the laser flash method [7] and thermal conductivity probes [8], are often inefficient. In theoretical calculations, solving the Boltzmann Transport Equation (BTE) based on Density Functional Theory (DFT) is regarded as the most reliable method for determining LTC [9, 10]. Molecular Dynamics (MD) offers an alternative [11]; however, the former is limited by its substantial computational resource requirements, while the latter’s accuracy depends on the choice of interatomic potentials [12]. In recent years, machine learning (ML) has emerged as a powerful and efficient data mining tool, gaining widespread application in materials science [13]. The intersection of materials science and artificial intelligence is commonly referred to as “Materials Informatics” [14] or the “Materials Genome” [15]. Early ML-based predictions of material properties primarily aimed at achieving high prediction accuracy [16, 17, 18]. Nevertheless, such efforts, relying on “black-box” models, offered limited support for advancing theoretical research in materials science. The focus has shifted towards model interpretability, prompting greater adoption of “white-box” models that contribute more substantially to theoretical advancements [19]. Research in materials informatics related to thermal conductivity has followed this trend. Efforts to predict LTC using black-box models such as Random Forest (RF) [20], Gaussian Process Regression (GPR) [21], and eXtreme Gradient Boosting (XGBoost) [22] have yielded reliable accuracy, whereas, the interpretability of these models is often hindered by their complexity. Alternatively, Genetic Programming-based Symbolic Regression (GPSR) [23] algorithms offer better interpretability, but their simplicity comes at the cost of reduced accuracy. The prevailing view is that the behavior and processes of complex models are difficult to understand and interpret, while simple models often lack strong fitting capabilities [24, 25, 19]. This presents a challenging trade-off, but the situation is gradually improving. The Sure Independence Screening and Sparsifying Operator (SISSO) [26], based on compressed sensing and symbolic regression, has not only surpassed the accuracy of the Slack semi-empirical model [27] but also narrowed the gap with black-box models like Kernel Ridge Regression (KRR) and GPR, and has been used to quantify feature sensitivity and identify key physical parameters influencing LTC [28]. Despite its potential, SISSO’s applicability is constrained by its high computational demands. Enumerating combinations of features and operators to construct descriptors poses an NP-hard problem [26], with resource requirements increasing for high-dimensional input features [29].
@@
Figure 1: A schematic illustration of the DL framework for LTC modeling and high-throughput prediction. This framework facilitates accurate and interpretable predictions of LTC for inorganic crystalline materials.
@@
Kolmogorov-Arnold Networks (KANs) [30] , a novel neural network architecture distinct from Multi-Layer Perceptrons (MLPs), show significant promise in solving partial differential equations (PDEs), time series prediction, and high-dimensional data tasks [31, 30, 32]. In some cases, smaller KAN models outperform MLPs [31, 30]. KANs have also demonstrated success in engineering applications, such as predicting chiller energy consumption [33] and nuclear binding energy [34], highlighting both feasibility and broad potential. Notably, their capabilities in symbolic regression offer interpretability advantages tabsent in MLPs [31, 30]. As previously mentioned, existing work has successfully used black-box models to accurately predict LTC. However, achieving accurate LTC predictions while simultaneously visualizing the model’s decision-making process, akin to white-box model, remains a significant challenge. The emergence of KANs provides an opportunity to address this issue.
@@
Figure 1 illustrates the framework of our work, which aims to model and predict LTC using the white-box deep learning model KAN and compare its performance with that of conventional black-box and white-box ML models. In addition to accuracy, we perform sensitivity analysis to examine interpretability differences across various white-box models and to assess the contributions of individual features to LTC. This approach further elucidates the relationships between LTC and key features, as well as feature-feature correlations. Additionally, Crystal Graph Convolutional Neural Network (CGCNN) [35] is employed to predict critical physical features required for LTC modeling with high accuracy, forming a two-stage framework that significantly reduces the time required for screening potential materials. This framework enables rapid qualitative assessments of thermal insulators and conductors. Finally, by integrating DFT and MD, our pretrained KAN model achieves outstanding accuracy in predicting the properties of candidates in an unlabeled extrapolation dataset.
@@ II Results
II.1 Establishment and evaluation of LTC models
@@
Figure 2: Model performance on dataset with complete features. (a) The blue circular markers represent samples from the training set, while the red star scatters are samples from the test set. (b) KAN model architecture based on the original dataset.
For regression models, the coefficient of determination ($R^{2}$) provides more information compared to other commonly used criteria [43]. As shown in Figure 2(a), on the original dataset, XGB slightly outperforms KAN, which in turn outperforms MLP; however, the differences are marginal, and all three models are obviously better than SISSO. Additionally, SISSO exhibits severe errors in a very limited subset of samples with extremely low LTC values (approximately $log(\kappa)<-1$). This phenomenon indicates that SISSO struggles with datasets characterized by class imbalance. SISSO tends to fit more to samples with a higher frequency and a narrower range of target values when dealing with imbalanced data, leading to suboptimal extrapolation performance. As a form of symbolic regression, SISSO inherently faces challenges in addressing sample imbalance. Common approaches to mitigate this issue in symbolic regression include resampling [44] and weighting [44, 45]. However, these methods inevitably impact the model’s interpretability [46, 47]. Therefore, we chose not to compromise SISSO’s interpretability solely for the purpose of enhancing its performance. It is noteworthy that XGBoost outperforms the deep learning models in this experiment, which can be attributed to several factors:
@@
$\log{(\kappa)_{{\rm{SISSO}}}}=-2.8+0.3\log\frac{{{\Theta^{2}}}}{{\left|{{C_{\rm{p}}}-{C_{\rm{v}}}}\right|}}-0.17\times{10^{7}}\gamma{K_{\rm{i}}}\sqrt[3]{{\Theta{S_{\rm{v}}}}}$ (1)
@@
the linear coefficients in the model are obtained through least-square regression. For the specific physical meanings of the features, please refer to Table 2. Our dataset consists of a total of 10 features, but the SISSO model utilizes only seven of these. Some features, such as the space group number, which are typically not considered to be related to LTC. SISSO effectively identifies and excludes such extraneous variables. However, it overlooks physical quantities that are theoretically linked to LTC, such as the thermal expansion coefficient [50, 51]. Our feature sensitivity analysis indicates that the thermal expansion coefficient is regarded as an important parameter for modeling LTC.
Table 2: Features adopted in the dataset.
Feature Abbreviation Explanation
space group relaxed $SG$ The space group number for the relaxed structure
AGL Debye temperature $\Theta$ Debye temperature
AGL Grüneisen parameter $\gamma$ Grüneisen parameter
AGL heat capacity constant volume (300K) ${{C}_{\text{v}}}$ Heat capacity per cell at constant volume (300K)
AGL heat capacity constant pressure (300K) ${{C}_{\text{p}}}$ Heat capacity per cell at constant pressure (300K)
AGL thermal expansion coefficient (300K) $\alpha$ Thermal expansion coefficient (300K)
AGL vibrational entropy atom (300K) ${{S}_{\text{v}}}$ Vibrational entropy per atom (300K)
AGL vibrational free energy atom (300K) $F$ Vibrational free energy per atom (300K)
AGL bulk modulus static (300K) ${{K}_{\text{s}}}$ Static bulk modulus (300K)
AGL bulk modulus isothermal (300K) ${{K}_{\text{i}}}$ Isothermal bulk modulus (300K)
@@
In this experiment, the network structure of the KAN model was also determined through automatic hyperparameter optimization via Optuna. The final architecture selected a dimensionality of 7 for the sum operation nodes (i.e., the network structure is $10\times 7\times 1$), with one KAN layer preceding and another following this operation. In the first KAN layer, which connects to the input layer, there are $10\times 7=70$ residual activation functions, while in the second KAN layer, connected to the output layer, the number is $7\times 1=7$. The detailed network architecture is shown in Figure 2(b). The MLP we designed features hidden layers with dimensions of $64\times 64\times 64$. This architecture balances robust pattern recognition capabilities and mitigating issues such as gradient explosion or vanishing gradients. Despite achieving comparable performance, KAN and MLP exhibit significant differences in terms of parameter count, with MLP generally requiring a substantially higher number of parameters. The analytical expressions provided by KAN are not equivalent to the model obtained through data training, unlike SISSO, where the resulting symbolic formula is exactly the model itself. In the KAN Layers, each residual activation function consists of a basis function and a linear combination of several B-spline functions. To symbolize the model, KAN selects the elementary functions with the highest linear correlation to the nodes as substitutes. After symbolization, KAN undergoes further training to determine the affine parameters for each symbolic function. Although this process enhances interpretability, it may slightly reduce performance. Nevertheless, in this experiment, the $R^{2}$ value of the symbolic formula derived from KAN, at 0.9655, remains higher than that obtained by SISSO. The symbolic model constructed by KAN is detailed in Supplementary Information Eq.( 1). The analytical expression derived by KAN is much more complex than that of SISSO, as each feature is involved in 7 operations, a consequence of the KAN architecture. Please note that each feature in this formula has undergone Min-Max normalization as part of the feature preprocessing. However, this preprocessing step does not affect the model’s interpretability, as it is easy to revert [52]. Despite the fact that the analytical expressions provided by KAN and SISSO accurately capture the relationship between features and LTC, there are a few points to consider:
@@
  • Generally, the symbolic models constructed through machine learning require more physical features (e.g., SISSO involves 7 features, while KAN utilizes all 10 features) compared to semi-empirical models (such as Slack [27], which, when the temperature variable is fixed at 300K, requires only 5 additional features). Moreover, obtaining the corresponding features for unknown materials often involves high experimental or computational costs, making such models appear less appealing at present.
@@
  • Both KAN and SISSO, as white-box models, achieve reliable predictive accuracy, but they differ in terms of interpretability. For instance, due to dimensional constraints (see Table S1), SISSO excludes certain features during model construction, and we cannot guarantee that the excluded features are necessarily redundant or useless. Similarly, in KAN, while different features have varying weights, we cannot ensure that all features genuinely influence LTC. Therefore, it is essential to introduce effective and robust quantitative methods to objectively evaluate the interpretability of these models.
@@ II.2 Deepening physical interpretability through sensitivity analysis
We identified the most critical features for LTC through two steps. The first step is correlation analysis, achieved by calculating correlation coefficients. The second step is sensitivity analysis, accomplished by computing sensitivity indices.
Figure 3: Feature correlation and sensitivity analysis. (a) A heatmap of correlation coefficients is presented, where the numbers in the upper-right triangle represent the Pearson correlation coefficients between pairs of features. In the lower-left triangle, the direction of the ellipse’s long axis indicates the positive or negative correlation, while the area reflects the strength of correlation. (b)-(e) Different sensitivity analysis methods are applied to evaluate feature importance for both the KAN and SISSO models.
When constructing the dataset, we curated a selection of closely related features, such as heat capacity at constant volume ($C_{v}$) and heat capacity at constant pressure ($C_{p}$), as well as isothermal bulk modulus ($K_{i}$) and static (adiabatic) bulk modulus ($K_{s}$). The purpose of this approach was to ensure that the ML models achieve reliable accuracy. The correlation coefficient heatmap can reveal the potential degree of linear association between features, as illustrated in Figure 3(a). It is important to clarify that $K_{i}$ and $K_{s}$ are not entirely identical concepts [53, 54, 55]. Both are temperature-dependent, and the difference between them increases as the temperature rises, but due to the minimal difference at 300K, a near-linear relationship appears, which is also observed between the two specific heat capacities [53]. Interestingly, in addition to the previously mentioned bulk moduli and heat capacities, the Debye temperature $\Theta$, vibrational free energy per atom $F$, and vibrational entropy per atom ${S_{\rm{v}}}$ also show a high degree of linear correlation. The quantitative relationship between vibrational free energy and vibrational entropy is defined by the following equation [56]:
@@
where $Q$ represents the total enthalpy of formation of the compound, denotes Kelvin temperature (here $T=300K$), and the equation above can be rewritten as $\frac{F}{{{S_{\rm{v}}}}}=\frac{Q}{{{S_{\rm{v}}}}}-T$. Due to the presence of the entropy-enthalpy compensation (EEC) [57] effect, a linear relationship exists between entropy and enthalpy, which is further reflected in the linear correlation between $S_{v}$ and $F$. Based on the work of G. D. Garbulsky and G. Ceder [58], within the harmonic approximation, when the temperature is greater than the system’s characteristic Debye temperature, the lattice Hamiltonian can be written as:
$H\left({\vec{\sigma},T}\right)={E_{0}}\left({\vec{\sigma}}\right)+\left\langle{\ln\left(\omega\right)}\right\rangle\left(\sigma\right){k_{\rm{B}}}T+\frac{{\hbar\left\langle{{\omega^{2}}}\right\rangle\left({\vec{\sigma}}\right)}}{{24{k_{\rm{B}}}T}}-\frac{{{\hbar^{4}}\left\langle{{\omega^{4}}}\right\rangle\left({\vec{\sigma}}\right)}}{{2880k_{\rm{B}}^{3}{T^{3}}}}+\cdots$ (3)
@@
The sensitivity analysis results in Figure 3(c) and (e) are relatively reasonable and satisfy the premise of ”similar sensitivities for correlated features,” but differences still exist. For instance, in Figure 3(e), the total effect indices of $C_{p}$ and $C_{v}$ are higher than those of other features, yet their first-order effect indices are approximately zero. A total effect index slightly higher than the first-order effect index indicates that the independent contributions of the features themselves are minimal, but their combinations with other features exert additional influence on the model output. In the symbolic SISSO model (Eq. 1), specific heat capacities appear in the form $\left|{{C_{v}}-{C_{p}}}\right|$. Such interaction terms reduce the contribution of individual specific heat variables but result in a higher total contribution. However, a first-order effect index of zero suggests that log($\kappa$) is entirely independent of specific heat capacities.
@@
When considering empirical models like Slack’s, the sensitivity analysis results of the SISSO symbolic model are contradictory and unsustainable, as they underestimate the influence of specific heat capacities on $log(\kappa)$. Furthermore, despite the presence of other highly correlated variables, the removal of redundant features causes a noticeable performance degradation in the SISSO model, as shown in Figure 4(a). This indicates that SISSO fails to accurately capture the correlations between physical quantities and exhibits poor robustness, whereas the KAN model shows the opposite behavior. For a primary feature, a ${K_{\rm{i}}}$ value exceeding ${{K}^{\text{T}}}$ indicates a strong correlation with other features, while ${{K}^{\text{T}}}=0$ signifies an exact correlation with other inputs [28, 66]. From this perspective, KAN is more effective than SISSO at overcoming feature dependency. This advantage can be attributed to KAN’s neural network architecture. As illustrated in Figure 4(b), in the absence of pruning, the neurons in KAN’s hidden layers are fully connected to both preceding and subsequent layers. This design allows each feature to independently and equitably contribute to log($\kappa$). Simultaneously, KAN leverages backpropagation to update the weights of various nodes, ensuring the model achieves optimal accuracy. This process not only guarantees fair participation of all features in model training but also effectively distinguishes their relative contribution.
@@
Figure 4: Performance comparison before and after dimensionality reduction. (a) After extracting key features and re-modeling the LTC, both KAN and the black-box models exhibited good robustness, while the robustness of SISSO did not meet expectations. (b) The simplified architecture of KAN model.
@@
Ultimately, we retained $C_{v}$, $\alpha$, $F$, and $K_{s}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of LTC with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq.13), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.

### v8

@@ I Introduction
Lattice thermal conductivity (LTC, $\kappa_{\rm L}$) is a critical physical parameter that quantifies a material’s ability to transfer heat through lattice vibrations. It has diverse applications in thermal management [1], energy conversion [2], and thermoelectric materials [3]. High-$\kappa_{\rm L}$ materials are used in electronic devices to efficiently dissipate heat and prevent overheating [4]. In contrast, low-$\kappa_{\rm L}$ materials exhibit excellent performance in thermoelectric conversion, making them ideal for developing efficient thermoelectric generators (TEGs) [5] and coolers [6]. $\kappa_{\rm L}$ is crucial for material design and optimization, yet acquiring it for specific materials is challenging.
@@
Traditional experimental methods for measuring $\kappa_{\rm L}$, such as the laser flash method [7] and thermal conductivity probes [8], are often inefficient. In theoretical calculations, solving the Boltzmann Transport Equation (BTE) based on Density Functional Theory (DFT) is regarded as the most reliable method for determining $\kappa_{\rm L}$ [9, 10]. Molecular Dynamics (MD) offers an alternative [11]; however, the former is limited by its substantial computational resource requirements, while the latter’s accuracy depends on the choice of interatomic potentials [12]. In recent years, machine learning (ML) has emerged as a powerful and efficient data mining tool, gaining widespread application in materials science [13]. The intersection of materials science and artificial intelligence is commonly referred to as “Materials Informatics” [14] or the “Materials Genome” [15]. Early ML-based predictions of material properties primarily aimed at achieving high prediction accuracy [16, 17, 18]. Nevertheless, such efforts, relying on “black-box” models, offered limited support for advancing theoretical research in materials science. The focus has shifted towards model interpretability, prompting greater adoption of “white-box” models that contribute more substantially to theoretical advancements [19]. Research in materials informatics related to thermal conductivity has followed this trend. Efforts to predict $\kappa_{\rm L}$ using black-box models such as Random Forest (RF) [20], Gaussian Process Regression (GPR) [21], and eXtreme Gradient Boosting (XGBoost) [22] have yielded reliable accuracy, whereas, the interpretability of these models is often hindered by their complexity. Alternatively, Genetic Programming-based Symbolic Regression (GPSR) [23] algorithms offer better interpretability, but their simplicity comes at the cost of reduced accuracy. The prevailing view is that the behavior and processes of complex models are difficult to understand and interpret, while simple models often lack strong fitting capabilities [24, 25, 19]. This presents a challenging trade-off, but the situation is gradually improving. The Sure Independence Screening and Sparsifying Operator (SISSO) [26], based on compressed sensing and symbolic regression, has not only surpassed the accuracy of the Slack semi-empirical model [27] but also narrowed the gap with black-box models like Kernel Ridge Regression (KRR) and GPR, and has been used to quantify feature sensitivity and identify key physical parameters influencing $\kappa_{\rm L}$ [28]. Despite its potential, SISSO’s applicability is constrained by its high computational demands. Enumerating combinations of features and operators to construct descriptors poses an NP-hard problem [26], with resource requirements increasing for high-dimensional input features [29].
@@
Figure 1: A schematic illustration of the DL framework for $\kappa_{\rm L}$ modeling and high-throughput prediction. This framework facilitates accurate and interpretable predictions of $\kappa_{\rm L}$ for inorganic crystalline materials.
@@
Kolmogorov-Arnold Networks (KANs) [30] , a novel neural network architecture distinct from Multi-Layer Perceptrons (MLPs), show significant promise in solving partial differential equations (PDEs), time series prediction, and high-dimensional data tasks [31, 30, 32]. In some cases, smaller KAN models outperform MLPs [31, 30]. KANs have also demonstrated success in engineering applications, such as predicting chiller energy consumption [33] and nuclear binding energy [34], highlighting both feasibility and broad potential. Notably, their capabilities in symbolic regression offer interpretability advantages that absent in MLPs [31, 30]. As previously mentioned, existing work has successfully used black-box models to accurately predict $\kappa_{\rm L}$. However, achieving accurate $\kappa_{\rm L}$ predictions while simultaneously visualizing the model’s decision-making process, akin to white-box model, remains a significant challenge. The emergence of KANs provides an opportunity to address this issue.
@@
Figure 1 illustrates the framework of our work, which aims to model and predict $\kappa_{\rm L}$ using the white-box deep learning model KAN and compare its performance with that of conventional black-box and white-box ML models. In addition to accuracy, we perform sensitivity analysis to examine interpretability differences across various white-box models and to assess the contributions of individual features to $\kappa_{\rm L}$. This approach further elucidates the relationships between $\kappa_{\rm L}$ and key features, as well as feature-feature correlations. Additionally, Crystal Graph Convolutional Neural Network (CGCNN) [35] is employed to predict critical physical features required for $\kappa_{\rm L}$ modeling with high accuracy, forming a two-stage framework that significantly reduces the time required for screening potential materials. This framework enables rapid qualitative assessments of thermal insulators and conductors. Finally, by integrating DFT and MD, our pretrained KAN model achieves outstanding accuracy in predicting the properties of candidates in an unlabeled extrapolation dataset.
@@ II Results
II.1 Establishment and evaluation of $\kappa_{\rm L}$ models
@@
Figure 2: Model performance on primary dataset. (a) The blue circular markers represent samples from the training set, while the red star scatters are samples from the test set. (b) KAN model architecture based on the primary dataset.
For regression models, the coefficient of determination ($R^{2}$) provides more information compared to other commonly used criteria [43]. As shown in Figure 2(a), on the primary dataset, XGB slightly outperforms KAN, which in turn outperforms MLP; however, the differences are marginal, and all three models are obviously better than SISSO. Additionally, SISSO exhibits severe errors in a very limited subset of samples with extremely low $\kappa_{\rm L}$ values (approximately $\log(\kappa_{\rm L})<-1$). This phenomenon indicates that SISSO struggles with datasets characterized by class imbalance. SISSO tends to fit more to samples with a higher frequency and a narrower range of target values when dealing with imbalanced data, leading to suboptimal extrapolation performance. As a form of symbolic regression, SISSO inherently faces challenges in addressing sample imbalance. Common approaches to mitigate this issue in symbolic regression include resampling [44] and weighting [44, 45]. However, these methods inevitably impact the model’s interpretability [46, 47]. Therefore, we chose not to compromise SISSO’s interpretability solely for the purpose of enhancing its performance. It is noteworthy that XGBoost outperforms the deep learning models in this experiment, which can be attributed to several factors:
@@
$\log{(\kappa)_{{\rm{SISSO}}}}=$ $-2.8+0.3\log\frac{{{\Theta^{2}}}}{{\left|{{C_{\rm{p}}}-{C_{\rm{v}}}}\right|}}$
$\quad-0.17\times{10^{7}}\gamma{K_{\rm{i}}}\sqrt[3]{{\Theta{S_{\rm{v}}}}}$ (1)
@@
the linear coefficients in the model are obtained through least-square regression. For the specific physical meanings of the features, please refer to Table 2. Our dataset consists of a total of 10 features, but the SISSO model utilizes only seven of these. Some features, such as the space group number, which are typically not considered to be related to $\kappa_{\rm L}$. SISSO effectively identifies and excludes such extraneous variables. However, it overlooks physical quantities that are theoretically linked to $\kappa_{\rm L}$, such as the thermal expansion coefficient [50, 51]. Our feature sensitivity analysis indicates that the thermal expansion coefficient is regarded as an important parameter for modeling $\kappa_{\rm L}$.
Table 2: Features adopted in the primary dataset.
Feature Abbreviation Unit Range
The space group number for the relaxed structure $SG$ Unitless -
Debye temperature $\Theta$ K [24.9909, 2144.93]
Grüneisen parameter $\gamma$ Unitless [0.00210114, 3.86399]
Heat capacity per cell at constant volume (300K) ${{C}_{\text{v}}}$ $k_{\rm B}$/cell [1.11566, 275.775]
Heat capacity per cell at constant pressure (300K) ${{C}_{\text{p}}}$ $k_{\rm B}$/cell [0.0, 286.301]
Thermal expansion coefficient (300K) $\alpha$ K-1 [$-3.17\times 10^{-6}$, $5.66601\times 10^{-4}$]
[…]
Vibrational free energy per atom (300K) $F$ meV/atom [$-233.268$, 210.145]
Static bulk modulus (300K) ${{K}_{\text{s}}}$ GPa [0.0, 440.158]
Isothermal bulk modulus (300K) ${{K}_{\text{i}}}$ GPa [0.0, 437.329]
@@
In this experiment, the network structure of the KAN model was also determined through automatic hyperparameter optimization via Optuna. The final architecture selected a dimensionality of 7 for the sum operation nodes (i.e., the network structure is $10\times 7\times 1$), with one KAN layer preceding and another following this operation. In the first KAN layer, which connects to the input layer, there are $10\times 7=70$ residual activation functions, while in the second KAN layer, connected to the output layer, the number is $7\times 1=7$. The detailed network architecture is shown in Figure 2(b). The MLP we designed features hidden layers with dimensions of $64\times 64\times 64$. This architecture balances robust pattern recognition capabilities and mitigating issues such as gradient explosion or vanishing gradients. Despite achieving comparable performance, KAN and MLP exhibit significant differences in terms of parameter count, with MLP generally requiring a substantially higher number of parameters. The analytical expressions provided by KAN are not equivalent to the model obtained through data training, unlike SISSO, where the resulting symbolic formula is exactly the model itself. In the KAN Layers, each residual activation function consists of a basis function and a linear combination of several B-spline functions. To symbolize the model, KAN selects the elementary functions with the highest linear correlation to the nodes as substitutes. After symbolization, KAN undergoes further training to determine the affine parameters for each symbolic function. Although this process enhances interpretability, it may slightly reduce performance. Nevertheless, in this experiment, the $R^{2}$ value of the symbolic formula derived from KAN, at 0.9655, remains higher than that obtained by SISSO. The symbolic model constructed by KAN is detailed in Supplementary Information Eq. (1). The analytical expression derived by KAN is much more complex than that of SISSO, as each feature is involved in 7 operations, a consequence of the KAN architecture. Please note that each feature in this formula has undergone Min-Max normalization as part of the feature preprocessing. However, this preprocessing step does not affect the model’s interpretability, as it is easy to revert [52]. Despite the fact that the analytical expressions provided by KAN and SISSO accurately capture the relationship between features and $\kappa_{\rm L}$, there are a few points to consider:
@@
  • The symbolic models constructed through machine learning require more physical features (e.g., SISSO involves 7 features, while KAN utilizes all 10 features) compared to semi-empirical models (such as Slack [27], which, when the temperature variable is fixed at 300K, requires only 5 additional features). Moreover, obtaining the corresponding features for unknown materials often involves high experimental or computational costs, making such models appear less appealing at present.
@@
  • Both KAN and SISSO achieve reliable predictive accuracy, but they differ in terms of interpretability. For instance, due to dimensional constraints (see Table S1), SISSO excludes certain features during model construction, and we cannot guarantee that the excluded features are necessarily redundant or useless. Similarly, in KAN, while different features have varying weights, we cannot ensure that all features genuinely influence $\kappa_{\rm L}$. Therefore, it is essential to introduce effective and robust quantitative methods to objectively evaluate the interpretability of these models.
@@ II.2 Deepening physical interpretability through sensitivity analysis
We identified the most critical features for $\kappa_{\rm L}$ through two steps. The first step is correlation analysis, achieved by calculating correlation coefficients. The second step is sensitivity analysis, accomplished by computing sensitivity indices.
Figure 3: Feature correlation and sensitivity analysis. (a) Heatmap of the correlation matrix. The numbers in each cell represent the Pearson correlation coefficients between the corresponding features on the horizontal and vertical axes. The color gradient visually reflects the pairwise correlation strength. (b)-(e) Different sensitivity analysis methods are applied to evaluate feature importance for both the KAN and SISSO models.
When constructing the dataset, we curated a selection of closely related features, such as heat capacity at constant volume ($C_{\rm v}$) and heat capacity at constant pressure ($C_{\rm p}$), as well as isothermal bulk modulus ($K_{\rm i}$) and static (adiabatic) bulk modulus ($K_{\rm s}$). The purpose of this approach was to ensure that the ML models achieve reliable accuracy. The correlation coefficient heatmap can reveal the potential degree of linear association between features, as illustrated in Figure 3(a). It is important to clarify that $K_{\rm i}$ and $K_{\rm s}$ are not entirely identical concepts [53, 54, 55]. Both are temperature-dependent, and the difference between them increases as the temperature rises, but due to the minimal difference at 300K, a near-linear relationship appears, which is also observed between the two specific heat capacities [53]. Interestingly, in addition to the previously mentioned bulk moduli and heat capacities, the Debye temperature $\Theta$, vibrational free energy per atom $F$, and vibrational entropy per atom ${S_{\rm{v}}}$ also show a high degree of linear correlation. The quantitative relationship between vibrational free energy and vibrational entropy is defined by the following equation [56]:
@@
where $Q$ represents the total enthalpy of formation of the compound, denotes Kelvin temperature (here $T=300{\rm K}$), and the equation above can be rewritten as $\frac{F}{{{S_{\rm{v}}}}}=\frac{Q}{{{S_{\rm{v}}}}}-T$. Due to the presence of the entropy-enthalpy compensation (EEC) [57] effect, a linear relationship exists between entropy and enthalpy, which is further reflected in the linear correlation between $S_{\rm v}$ and $F$. Based on the work of G. D. Garbulsky and G. Ceder [58], within the harmonic approximation, when the temperature is greater than the system’s characteristic Debye temperature, the lattice Hamiltonian can be written as:
$H\left({\vec{\sigma},T}\right)=$ ${E_{0}}\left({\vec{\sigma}}\right)+\left\langle{\ln\left(\omega\right)}\right\rangle\left(\sigma\right){k_{\rm{B}}}T$
$+\frac{{\hbar\left\langle{{\omega^{2}}}\right\rangle\left({\vec{\sigma}}\right)}}{{24{k_{\rm{B}}}T}}-\frac{{{\hbar^{4}}\left\langle{{\omega^{4}}}\right\rangle\left({\vec{\sigma}}\right)}}{{2880k_{\rm{B}}^{3}{T^{3}}}}+\cdots$ (3)
@@
The sensitivity analysis results in Figure 3(c) and (e) are relatively reasonable and satisfy the premise of ”similar sensitivities for correlated features,” but differences still exist. For instance, in Figure 3(e), the total effect indices of $C_{\rm p}$ and $C_{\rm v}$ are higher than those of other features, yet their first-order effect indices are approximately zero. A total effect index slightly higher than the first-order effect index indicates that the independent contributions of the features themselves are minimal, but their combinations with other features exert additional influence on the model output. In the symbolic SISSO model (Eq. (1)), specific heat capacities appear in the form $\left|{{C_{\rm v}}-{C_{\rm p}}}\right|$. Such interaction terms reduce the contribution of individual specific heat variables but result in a higher total contribution. However, a first-order effect index of zero suggests that $\log(\kappa_{\rm L})$ is entirely independent of specific heat capacities.
@@
When considering empirical models like Slack’s, the sensitivity analysis results of the SISSO symbolic model are contradictory and unsustainable, as they underestimate the influence of specific heat capacities on $\log(\kappa_{\rm L})$. Furthermore, despite the presence of other highly correlated variables, the removal of redundant features causes a noticeable performance degradation in the SISSO model, as shown in Figure 4(a). This indicates that SISSO fails to accurately capture the correlations between physical quantities and exhibits poor robustness, whereas the KAN model shows the opposite behavior. For a primary feature, a ${K_{\rm{i}}}$ value exceeding ${{K}^{\text{T}}}$ indicates a strong correlation with other features, while ${{K}^{\text{T}}}=0$ signifies an exact correlation with other inputs [28, 66]. From this perspective, KAN is more effective than SISSO at overcoming feature dependency. This advantage can be attributed to KAN’s neural network architecture. As illustrated in Figure 4(b), in the absence of pruning, the neurons in KAN’s hidden layers are fully connected to both preceding and subsequent layers. This design allows each feature to independently and equitably contribute to $\log(\kappa_{\rm L})$. Simultaneously, KAN leverages backpropagation to update the weights of various nodes, ensuring the model achieves optimal accuracy. This process not only guarantees fair participation of all features in model training but also effectively distinguishes their relative contribution.
@@
Figure 4: Performance comparison before and after dimensionality reduction. (a) After extracting key features and re-modeling the $\kappa_{\rm L}$, both KAN and the black-box models exhibited good robustness, while the robustness of SISSO did not meet expectations. (b) The simplified architecture of KAN model.
@@
Ultimately, we retained $C_{\rm v}$, $\alpha$, $F$, and $K_{\rm s}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of $\kappa_{\rm L}$ with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq. (12)), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.